OpenStudy (anonymous):

Will Become A Fan and Give Medal/Testimonial Devon wants to write an equation for a line that passes through 2 of the data points he has collected. The points are (8, 5) and (–12, –9). He writes the equation 7x – 10y = 3. Is this a good model? Explain your reasoning.

OpenStudy (anonymous):

@ganeshie8 @zepdrix

OpenStudy (studygurl14):

Use slope formula to find slope \[\frac{ -9-5 }{ -12-8 }=\frac{ -14 }{ -20 }=\frac{ 7 }{ 10 }\] 7/10 = m = slope y= mx + b y= (7/10)x + b (8,5) (5) = (7/10)(8) + b b = -3/5 So.... \[y=\frac{ 7 }{ 10 }x-\frac{ 3 }{ 5 }\] multiply by 10 \[10y = 7x -6\] --> -7x + 10y = -6 --> 7x - 10y = 6 So, if 7x - 10y = 3 a good model?